# GalaxyGenius: A Mock Galaxy Image Generator for Various Telescopes from Hydrodynamical Simulations

### v1

@@ Abstract
Aims. To advance research on galaxies in the era of large-scale sky surveys, we introduce GalaxyGenius, a Python package designed to produce synthetic galaxy images tailored to different telescopes based on hydrodynamical simulations.
Methods. The package comprises three main modules: data preprocessing, ideal data cube generation, and mock observation. Specifically, the preprocessing module extracts necessary properties of star and gas particles for a selected subhalo from hydrodynamical simulations and creates the execution file for the following radiative transfer procedure. Subsequently, building on the above information, the ideal data cube generation module executes a widely-used radiative transfer project, specifically the SKIRT, to perform the SED assignment for each particle and the radiative transfer procedure to produce an IFU-like ideal data cube. Lastly, the mock observation module takes the ideal data cube and applies the throughputs of aiming telescopes while also incorporating the relevant instrumental effects, point spread functions (PSFs), and background noise to generate desired mock observational images of galaxies.
Results. To showcase the outcomes of GalaxyGenius, We create a series of mock images of galaxies based on the IllustrisTNG and EAGLE simulations for both space and ground-based surveys spanning from the ultraviolet to infrared wavelength coverages, including CSST, Euclid, HST, JWST, Roman, and HSC.
Conclusions. GalaxyGenius offers a flexible framework developed to generate mock galaxy images with customizable recipes. These generated images can serve as valuable references for verifying and validating new approaches in astronomical research, and they can also act as training sets for deep learning relevant studies where real observational data are insufficient.
@@
Galaxies are fundamental building blocks of the Universe, serving as cosmic laboratories where stars, planetary systems, and various astrophysical phenomena are born and evolved. In astronomy, the study of galaxies provides insights into the processes of star formation, stellar dynamics, and the interstellar medium, enhancing our understanding of the lifecycle of celestial objects (Madau & Dickinson, 2014; Conselice, 2014; Behroozi et al., 2013; Kravtsov & Borgani, 2012). In cosmology, galaxies are crucial for mapping the large-scale structure of the Universe, offering clues on the distribution of dark matter and the influence of dark energy on cosmic expansion (Wang et al., 2020; Wang & Zhao, 2020; Springel et al., 2006; Weinberg et al., 2013). Observations of galaxies at different redshift allow scientists to look back in time and study the Universe’s evolution, shedding light on the formation and growth of cosmic structures over billions of years.
@@
Currently, the most sophisticated and realistic galaxies are simulated or generated from hydrodynamical simulations. These categories of simulations are distinct from N-body simulations which only consider the evolution of dark matter. Instead, hydrodynamical simulations are computational tools that model the behavior of fluids, such as gases and liquids, under various physical conditions. In astrophysics, hydrodynamical simulations are crucial for understanding complex phenomena like galaxy formation, star formation, and the dynamics of interstellar and intergalactic media (Nelson et al., 2019b; Genel et al., 2014). By solving the equations of fluid dynamics, often coupled with additional physics like gravity, magnetic fields, gas cooling and feedback from supernovae and active galactic nuclei (AGN) and radiation, researchers can recreate and study the evolution of cosmic structures over time and reproduce many observed properties of galaxies (Crain & van de Voort, 2023; Valentini & Dolag, 2025). They help elucidate the role of feedback mechanisms in regulating star formation and shaping the interstellar medium. Additionally, by comparing simulation results with observational data, researchers can test and refine theoretical models, leading to a more comprehensive understanding of the universe’s history and structure. Utill now, hydrodynamical simulations, such as Illustris (Vogelsberger et al., 2014; Genel et al., 2014; Sijacki et al., 2015), IllustrisTNG (Marinacci et al., 2018; Pillepich et al., 2018, 2019; Nelson et al., 2018, 2019a; Naiman et al., 2018; Springel et al., 2018), EAGLE (The EAGLE team, 2017; McAlpine et al., 2016), FLAMINGO (Schaye et al., 2023), and other simulations, are widely utilized to address the challenges in galaxy formation and evolution and promote our understanding of our Universe.
@@
From hydrodynamical simulations, galaxies can be generated by assigning particles in a subhalo spectral energy distributions (SEDs) based on particle properties. To account for the dust effects, radiative transfer processes or simplified prescriptions shown in Martin et al. (2022); Fortuni et al. (2023); LaChance et al. (2025) can be conducted to model the absorption, scattering and emissions. This process typically generates 3-dimensional ideal data cubes. Subsequently, the data cubes are convolved with throughputs of filters for specific survey to obtain the bandpass images. Finally, point spread function (PSF) and instrumental effects are applied to generate mock galaxy images. Several endeavors exist for generating galaxies from the EAGLE (Crain et al., 2015; Schaye et al., 2015) and Illustris (Vogelsberger et al., 2014; Genel et al., 2014; Sijacki et al., 2015) simulations, as well as its upgraded version, IllustrisTNG (Marinacci et al., 2018; Pillepich et al., 2018, 2019; Nelson et al., 2018, 2019a; Naiman et al., 2018; Springel et al., 2018), employing radiative transfer codes, such as DIRTY (Gordon et al., 2001), SKIRT (Baes et al., 2011; Camps & Baes, 2015), HYPERION (Robitaille, 2011), RADMC-3D (Dullemond et al., 2012), POLARIS (Reissl et al., 2016), and SOC (Juvela, 2019).
@@
These generated galaxies serve as a versatile platform for extensive applications in galaxy surveys. One such application is machine learning-based studies, particularly those utilizing deep learning techniques, such as estimation of photometric redshifts (Naidoo et al., 2023; Moskowitz et al., 2024; Zhou et al., 2022a, b), classification of galaxy morphologies (Pfeffer et al., 2023; Hambleton et al., 2011; Dickinson et al., 2018), measurement of galaxy properties (shear, stellar mass) (Euclid Collaboration et al., 2024b; Tewes et al., 2019; Zhang et al., 2024; Bonjean et al., 2019; Chu et al., 2024), identification of galaxy mergers (Ferreira et al., 2020; Pearson et al., 2019; Nevin et al., 2019; Snyder et al., 2019; Bottrell et al., 2019; Guzmán-Ortega et al., 2023; Rose et al., 2023; Margalef-Bentabol et al., 2024; Wilkinson et al., 2024), and investigation of blending effects and deblending algorithms (Arcelin et al., 2021; Burke et al., 2019; Hemmati et al., 2022). These galaxies can effectively address the insufficiency of training data in these studies, and can aid in the development of a robust deep learning model. Moreover, these galaxies provide a controlled and well-defined environment where ground-truth information is precisely known. This allows us to test, validate and calibrate the performance of deep learning models under various observational conditions, such as noise, resolution, and PSF effects. Furthermore, by generating mock observations that mimic real survey data, we can assess how well models generalize to real-world datasets.
@@
Several projects are dedicated to generating mock observational galaxy images from hydrodynamical simulations, such as FORECAST (Fortuni et al., 2023) and RealSim (Bottrell et al., 2017, 2019). These projects employ simplified prescriptions, including semi-analytic models (Guiderdoni & Rocca-Volmerange, 1987; Devriendt & Guiderdoni, 2000; Nelson et al., 2019b; Vogelsberger et al., 2020), to account for absorption, scattering, and re-emission of dust components. While in our work, we present a Python package, GalaxyGenius, that is capable of generating galaxies in ultraviolet to infrared bands, incorporating comprehensive radiative transfer procedures implemented by SKIRT. Although the radiative transfer process is computationally intensive, this approach generates galaxies with enhanced physical details and realism, as it self-consistently models the interplay between radiation and interstellar dust. This package integrates three modules: data preprocessing, ideal data cube generation, and mock observation. The preprocessing encompasses extraction of particles and corresponding properties used for SED assignment from hydrodynamical simulations, and creation of execution file for radiative transfer code. Subsequently, data cube generation module produces an IFU-like data by executing SKIRT based on information obtained from preprocessing. Finally, mock observation module generate realistic observational images by incorporating instrumental effects specific for photometric instruments. We demonstrate the capability of GalaxyGenius by creating mock images from IllustrisTNG and EAGLE simulations for both space and ground-based surveys spanning from ultraviolet to infrared wavelength coverages, including the Chinese Space Station Telescope (CSST, Zhan (2018); Gong et al. (2019)), Hubble Space Telescope (HST, Freedman et al. (2001); Momcheva et al. (2016)), James Webb Space Telescope (JWST, Sabelhaus & Decker (2004)), Euclid Space Telescope (Euclid, Euclid Collaboration et al. (2024d)), Nancy Grace Roman Space Telescope (Roman, Spergel et al. (2015)), and Hyper Suprime-Cam (HSC, Aihara et al. (2018)). Additionally, we also investigate the effects of dust elements on galaxy generation, particularly for dust recipes and dust models. Furthermore, we discuss some limitations on generating galaxies from hydrodynamical simulations using SKIRT project.
@@
This paper is organized as follows: Section 2 elaborates the implementation details of GalaxyGenius on generating galaxies from hydrodynamical simulations using SKIRT. We demonstrate some sample outputs for various photometric surveys in Section 3. Subsequently, in Section 4, we discuss the effects of dust recipes and dust models, and explain some limitations of generating galaxies from hydrodynamical simulations. Finally, this work is summarized in Section 5. Appendix A describe the instrumental parameters for each considered survey and their origins. Appendix B displays PSFs utilized for photometric bands for each considered survey. Appendix C provides the derivation details for noise levels achieved from limiting magnitudes.
@@ 2.1 Overall framework
The overall framework of GalaxyGenius is illustrated in Figure 1, which comprises three primary modules for generating mock observational galaxy images tailored to specific surveys:
(i) Data preprocessing: This module extracts coordinates and relevant properties for different categories of particles, including stars, star-forming regions and dust elements, for a selected subhalo of hydrodynamical simulation. Additionally, it also prepares the execution file for radiative transfer codes based on the particle files and configurations. For further details, please refer to Section 2.2.
(ii) Ideal data cube generation: This module executes the radiative transfer procedures based on the particle files and execution file created in data preprocessing. Finally, an IFU-like ideal data cubes and SEDs for different viewing angles can be obtained. Please refer to Section 2.3 for more details.
(iii) Mock observation: This module aims to construct mock galaxy images anticipated for specific survey. It initially creates ideal bandpass images from the data cube outputted above, utilizing the throughput of each filter. Subsequently, it applies instrumental effects, including point spread functions (PSF), background and instrumental noises under specified observational conditions. The details are described in Section 2.4.
@@ 2.2 Data preprocessing
SFR | Calculated | Star-formation rate of the HII region |
@@
Here we take snapshot-94 () of TNG100 from IllustrisTNG as a case study. Firstly, subhalos are read from the subhalo catalog derived using the Subfind algorithm (Springel et al., 2000). The subhalos in desired stellar mass range should have a constraint as SubhaloFlag == 1, which is to ensure that these subhalos can be considered as galaxy or satellite of cosmological origin (de los Rios et al., 2023; Nelson et al., 2019b).
@@
We follow the procedure established by Rodriguez-Gomez et al. (2019). For each particle older than 10 Myr, stellar population SED from the Bruzual & Charlot (BC03, Bruzual & Charlot (2003)) or Flexible Stellar Population Synthesis (FSPS, Conroy & Gunn (2010)) can be employed. Here BC03 SSP family with Chabrier initial mass function (Chabrier, 2003) is assigned. We call them as stars or BC03 particles interchangeably in following discussions. The initial mass , metallicity , and age for the SSP are directly inherited from TNG data.
@@
While for particles younger than 10 Myr, it is assumed that they are still partially embedded within their respective clouds. Consequently, these particles are assigned SEDs based on a library of HII region templates derived from the MAPPINGS-III library (Groves et al., 2008). We call these particles as star-forming regions in the following context. Each particle’s SED is characterized by five parameters. Among these parameters, metallicity can be directly retrieved from particle data. The SFR is assumed to remain constant throughout HII region’s lifetime of 10 Myr. On the other hand, the remaining three parameters are free ones. The compactness determines the dust temperature distribution and, consequently, influences the shape of far-infrared continuum dust emission. We sample compactness using a lognormal Gaussian distribution with a mean of and a standard deviation of 0.4 for each particle (Kapoor et al., 2021; Trčka et al., 2022). The interstellar medium (ISM) pressure only affects the strength of several emission lines but does not alter the shape of the SED (Groves et al., 2008). Therefore, we treat it as a constant value of . The final one, covering factor of the photodissociation region (PDR) describes the extent that HII regions are coated (Kaufman et al., 1999; Pound & Wolfire, 2022). Since the coating will eventually clear in PDR, this parameter is set as , where represents the age of the stellar particle, and is the molecular cloud clearing or dissipation time scale. means that this star-forming region is thoroughly transparent, while indicates that it is completely covered. We adopt the value of 3 Myr, as recommended in Trčka et al. (2022).
@@
The next step involves modeling the distribution of dust elements within each galaxy. The density of dust at any given volume is determined by the assumption that a predetermined percentage of the metals present in the interstellar medium (ISM) gas are encapsulated within dust grains, as follows:
@@
where and represents the metallicity and density of the gas respectively. And denotes the dust-to-metal ratio. This ratio is correlated with stellar mass, according to work by Lu et al. (2023), however, we assume it to be a constant of 0.3 for simplicity. Furthermore, we distinguish between the hot circumgalactic medium and the cooler ISM gas using two distinct dust recipes:
@@
(i) Only gas cells with non-zero SFR or with a temperature , below are considered as ISM (Camps et al., 2016):
@@
where is calculated from internal energy and electron abundance as recommended in TNG website 111https://www.tng-project.org/data/docs/faq/#gen6 as follows:
@@
where , , and are the adiabatic index, the Bolzmann constant, the hydrogen mass fraction and the mass of proton respectively. The condition captures gas particles that are eligible for star formation but were not actually converted into a star-forming region. Meanwhile, the temperature cutoff value of for non-star-forming gas particles ensures that only gas capable of forming dust is considered, as dust cannot form or is rapidly destroyed in hot gas ((Camps et al., 2016; Guhathakurta & Draine, 1989)). The cutoff temperature is arbitrary and empirical. However, this condition does not cause a substantial difference in the dust geometry (see Section 4.1), since the diffuse dust content of each galaxy is essentially determined by the star-forming gas alone (Rodriguez-Gomez et al., 2019; Kapoor et al., 2021).
@@
(ii) The rotationally supported interstellar gas, settled in the disc, are distinguished from the hot circumgalactic gas following Torrey et al. (2012):
@@
where is temperature calculated using Equation 3, and is gas density directly obtained from snapshot data. This condition has been shown to effectively remove cells in the hot halo (Torrey et al., 2012; McKinnon et al., 2016). To enhance the adaptability of GalaxyGenius, we implement both dust recipes in GalaxyGenius, allowing users to freely choose between them. Furthermore, in TNG, hydrodynamics for the gas component is solved using a moving Voronoi mesh via the AREPO code (Weinberger et al., 2020), eliminating the need for smoothing length. The necessary properties for stars, star-forming regions, and dust elements along with their origins are summarized in Table 1.
@@
We employ SKIRT to perform radiative transfer procedures. SKIRT is a versatile and widely used Monte Carlo radiative transfer code 222https://skirt.ugent.be/ designed to model the interaction of light with matter in astrophysical systems (Baes et al., 2011; Camps & Baes, 2015). Developed to study processes such as dust absorption, scattering, and thermal re-emission, SKIRT is capable of generating realistic synthetic observations of galaxies, star-forming regions, and other celestial objects. A key feature of SKIRT is its ability to incorporate complex geometries, spatial distributions of stars and dust, and wavelength-dependent properties of materials. Through Monte-Carlo techniques, SKIRT simulates the paths of photon packets as they propagate through a medium, accounting for scattering, absorption, and re-emission processes. SKIRT is particularly well-suited for creating mock observations from hydrodynamical simulations, such as IllustrisTNG and EAGLE (Nelson et al., 2018; Rodriguez-Gomez et al., 2019; Shen et al., 2020; Camps et al., 2016; Guzmán-Ortega et al., 2023; de los Rios et al., 2023; Bottrell et al., 2024; Baes et al., 2024). In such scenario, SKIRT transforms theoretical outputs, such as stellar population distributions and gas content, into observable quantities, including images and SEDs across various wavelength coverage, from ultraviolet to far-infrared. The execution file for SKIRT, with a extension, can be generated through a user-friendly command-line Q&A session.

### v2

# GalaxyGenius: Mock galaxy image generator for various telescopes from hydrodynamical simulations

@@ Abstract
Aims. We introduce GalaxyGenius, a Python package designed to produce synthetic galaxy images tailored to different telescopes based on hydrodynamical simulations. Its implementation will support and advance research on galaxies in the era of large-scale sky surveys,
Methods. The package comprises three main modules: data preprocessing, ideal data cube generation, and mock observation. Specifically, the preprocessing module extracts necessary properties of star and gas particles for a selected subhalo from hydrodynamical simulations and creates the execution file for the following radiative transfer procedure. Subsequently, building on the above information, the ideal data cube generation module executes a widely used radiative transfer project, specifically the SKIRT, to perform the SED assignment for each particle and the radiative transfer procedure to produce an IFU-like ideal data cube. Lastly, the mock observation module takes the ideal data cube and applies the throughputs of aiming telescopes, while also incorporating the relevant instrumental effects, point spread functions (PSFs), and background noise to generate the required mock observational images of galaxies.
Results. To showcase the outcomes of GalaxyGenius, we created a series of mock images of galaxies based on the IllustrisTNG and EAGLE simulations for both space and ground-based surveys, spanning ultraviolet (UV) to infrared (IR) wavelength coverage, including CSST, Euclid, HST, JWST, Roman, and HSC.
Conclusions. GalaxyGenius offers a flexible framework to generate mock galaxy images with customizable recipes. These generated images can serve as valuable references for verifying and validating new approaches in astronomical research. They can also serve as training sets for relevant studies using deep learning in cases where real observational data are insufficient.
@@
Galaxies are the fundamental building blocks of the Universe, serving as cosmic laboratories where stars, planetary systems, and various astrophysical phenomena are born and evolve. In astronomy, the study of galaxies provides insights into the processes of star formation, stellar dynamics, and the interstellar medium (ISM), enhancing our understanding of the life cycle of celestial objects (Madau & Dickinson, 2014; Conselice, 2014; Behroozi et al., 2013; Kravtsov & Borgani, 2012). In cosmology, galaxies are crucial for mapping the large-scale structure of the Universe, offering clues on the distribution of dark matter and the influence of dark energy on cosmic expansion (Wang et al., 2020; Wang & Zhao, 2020; Springel et al., 2006; Weinberg et al., 2013). Observations of galaxies at different redshift allow scientists to look back in time and study the Universe’s evolution, shedding light on the formation and growth of cosmic structures over billions of years.
@@
Currently, the most sophisticated and realistic galaxies are simulated or generated from hydrodynamical simulations. These categories of simulations are distinct from N-body simulations, which only consider the evolution of dark matter. Instead, hydrodynamical simulations are computational tools that model the behavior of fluids, such as gases and liquids, under various physical conditions. In astrophysics, hydrodynamical simulations are crucial for understanding complex phenomena like galaxy formation, star formation, and the dynamics of interstellar and intergalactic media (Nelson et al., 2019b; Genel et al., 2014). By solving the equations of fluid dynamics, often coupled with additional physics like gravity, magnetic fields, gas cooling and feedback from supernovae and active galactic nuclei (AGNs) and radiation, researchers can recreate and study the evolution of cosmic structures over time and reproduce many observed properties of galaxies (Crain & van de Voort, 2023; Valentini & Dolag, 2025). They help elucidate the role of feedback mechanisms in regulating star formation and shaping the ISM. Additionally, by comparing simulation results with observational data, researchers can test and refine theoretical models, leading to a more comprehensive understanding of the universe’s history and structure. Utill now, hydrodynamical simulations, such as Illustris (Vogelsberger et al., 2014; Genel et al., 2014; Sijacki et al., 2015), IllustrisTNG (Marinacci et al., 2018; Pillepich et al., 2018, 2019; Nelson et al., 2018, 2019a; Naiman et al., 2018; Springel et al., 2018), EAGLE (The EAGLE team, 2017; McAlpine et al., 2016), FLAMINGO (Schaye et al., 2023), and other simulations, are widely utilized to address the challenges in galaxy formation and evolution and promote our understanding of our Universe.
@@
From hydrodynamical simulations, galaxies can be generated by assigning particles in a subhalo spectral energy distributions (SEDs) based on particle properties. To account for the dust effects, radiative transfer processes or simplified prescriptions shown in Martin et al. (2022); Fortuni et al. (2023); LaChance et al. (2025) can be conducted to model the absorption, scattering, and emissions. This process typically generates three-dimensional (3D) ideal data cubes. Subsequently, the data cubes are convolved with throughputs of filters for specific survey to obtain the bandpass images. Finally, point spread function (PSF) and instrumental effects are applied to generate mock galaxy images. Several endeavors exist for generating galaxies from the EAGLE (Crain et al., 2015; Schaye et al., 2015) and Illustris (Vogelsberger et al., 2014; Genel et al., 2014; Sijacki et al., 2015) simulations, as well as its upgraded version, IllustrisTNG (Marinacci et al., 2018; Pillepich et al., 2018, 2019; Nelson et al., 2018, 2019a; Naiman et al., 2018; Springel et al., 2018), employing radiative transfer codes, such as DIRTY (Gordon et al., 2001), SKIRT (Baes et al., 2011; Camps & Baes, 2015), HYPERION (Robitaille, 2011), RADMC-3D (Dullemond et al., 2012), POLARIS (Reissl et al., 2016), and SOC (Juvela, 2019).
@@
These generated galaxies serve as a versatile platform for extensive applications in galaxy surveys. One such application is machine learning-based studies, particularly those utilizing deep learning techniques, such as the estimation of photometric redshifts (Naidoo et al., 2023; Moskowitz et al., 2024; Zhou et al., 2022a, b), classification of galaxy morphologies (Pfeffer et al., 2023; Hambleton et al., 2011; Dickinson et al., 2018; Gong et al., 2025), measurement of galaxy properties (shear, stellar mass) (Euclid Collaboration et al., 2024b; Tewes et al., 2019; Zhang et al., 2024; Bonjean et al., 2019; Chu et al., 2024; Tang et al., 2018, 2020, 2021, 2023), identification of galaxy mergers (Ferreira et al., 2020; Pearson et al., 2019; Nevin et al., 2019; Snyder et al., 2019; Bottrell et al., 2019; Guzmán-Ortega et al., 2023; Rose et al., 2023; Margalef-Bentabol et al., 2024; Wilkinson et al., 2024), and investigation of blending effects and deblending algorithms (Arcelin et al., 2021; Burke et al., 2019; Hemmati et al., 2022). These galaxies can effectively address the insufficiency of training data in these studies and can aid in the development of a robust deep learning model. Moreover, these galaxies provide a controlled and well-defined environment where ground-truth information is precisely known. This allows us to test, validate, and calibrate the performance of deep learning models under various observational conditions, such as noise, resolution, and PSF effects. Furthermore, by generating mock observations that mimic real survey data, we can assess how well models generalize to real-world datasets.
@@
Several projects are dedicated to generating mock observational galaxy images from hydrodynamical simulations, such as FORECAST (Fortuni et al., 2023) and RealSim (Bottrell et al., 2017, 2019). These projects employed simplified prescriptions, including semi-analytic models (Guiderdoni & Rocca-Volmerange, 1987; Devriendt & Guiderdoni, 2000; Nelson et al., 2019b; Vogelsberger et al., 2020), to account for absorption, scattering, and re-emission of dust components. While in our work, we present a Python package, GalaxyGenius, that is capable of generating galaxies in ultraviolet (UV) to infrared (IR) bands, incorporating comprehensive radiative transfer procedures implemented by SKIRT. Although the radiative transfer process is computationally intensive, this approach generates galaxies with enhanced physical details and realism, as it self-consistently models the interplay between radiation and interstellar dust. This package integrates three modules: data preprocessing, ideal data cube generation, and mock observation. The preprocessing encompasses extraction of particles and corresponding properties used for SED assignment from hydrodynamical simulations and creation of an execution file for the radiative transfer code. Subsequently, data cube generation module produces an IFU-like data by executing SKIRT based on information obtained from preprocessing. Finally, mock observation module generate realistic observational images by incorporating instrumental effects specific for photometric instruments. We demonstrate the capability of GalaxyGenius by creating mock images from IllustrisTNG and EAGLE simulations for both space and ground-based surveys spanning UV to IR wavelength coverage, including the Chinese Space Station Telescope (CSST, Zhan (2018); Gong et al. (2019)), Hubble Space Telescope (HST, Freedman et al. (2001); Momcheva et al. (2016)), James Webb Space Telescope (JWST, Sabelhaus & Decker (2004)), Euclid Space Telescope (Euclid, Euclid Collaboration et al. (2024d)), Nancy Grace Roman Space Telescope (Roman, Spergel et al. (2015)), and Hyper Suprime-Cam (HSC, Aihara et al. (2018)). Additionally, we also investigate the effects of dust elements on galaxy generation, particularly for dust recipes and dust models. Furthermore, we discuss some limitations on generating galaxies from hydrodynamical simulations using SKIRT project.
@@
This paper is organized as follows. Section 2 elaborates the implementation details of GalaxyGenius on generating galaxies from hydrodynamical simulations using SKIRT. We demonstrate some sample outputs for various photometric surveys in Section 3. Subsequently, in Section 4, we discuss the effects of dust recipes and dust models and explain some limitations of generating galaxies from hydrodynamical simulations. Finally, this work is summarized in Section 5. Appendix A describe the instrumental parameters for each considered survey and their origins. Appendix B displays PSFs utilized for photometric bands for each considered survey. Appendix C provides the derivation details for noise levels achieved from limiting magnitudes.
@@ 2.1 Overall framework
The overall framework of GalaxyGenius is illustrated in Figure 1, which comprises three primary modules for generating mock observational galaxy images tailored to specific surveys, described below.
(i) Data preprocessing: this module extracts coordinates and relevant properties for different categories of particles, including stars, star-forming regions and dust elements, for a selected subhalo of hydrodynamical simulation. Additionally, it also prepares the execution file for radiative transfer codes based on the particle files and configurations. For further details, we refer to Section 2.2.
(ii) Ideal data cube generation: this module executes the radiative transfer procedures based on the particle files and execution file created in data preprocessing. Finally, an IFU-like ideal data cubes and SEDs for different viewing angles can be obtained. We refer to Section 2.3 for more details.
(iii) Mock observation: this module aims to construct mock galaxy images anticipated for specific survey. It initially creates ideal bandpass images from the data cube outputted above, utilizing the throughput of each filter. Subsequently, it applies instrumental effects, including point spread functions (PSF) and background and instrumental noises under specified observational conditions. The details are described in Section 2.4.
@@ 2.2 Data preprocessing
SFR | Calculated | SFR of the HII region |
@@
Here, we take snapshot-94 () of TNG100 from IllustrisTNG as a case study. Firstly, subhalos are read from the subhalo catalog derived using the Subfind algorithm (Springel et al., 2000). The subhalos in the desired stellar mass range should have a constraint as SubhaloFlag == 1, which is to ensure that these subhalos can be considered as galaxy or satellite of cosmological origin (de los Rios et al., 2023; Nelson et al., 2019b).
@@
We follow the procedure established by Rodriguez-Gomez et al. (2019). For each particle older than 10 Myr, stellar population SED from the Bruzual & Charlot (BC03, Bruzual & Charlot (2003)) or Flexible Stellar Population Synthesis (FSPS, Conroy & Gunn (2010)) can be employed. Here, BC03 SSP family with Chabrier initial mass function (Chabrier, 2003) is assigned. We call them as stars or BC03 particles interchangeably in following discussions. The initial mass, , metallicity, , and age, , for the SSP are directly inherited from TNG data.
@@
While for particles younger than 10 Myr, it is assumed that they are still partially embedded within their respective clouds. Consequently, these particles are assigned SEDs based on a library of HII region templates derived from the MAPPINGS-III library (Groves et al., 2008). We call these particles as star-forming regions in the following context. Each particle’s SED is characterized by five parameters. Among these parameters, metallicity, , can be directly retrieved from particle data. The SFR is assumed to remain constant throughout HII region’s lifetime of 10 Myr. On the other hand, the remaining three parameters are free ones. The compactness, , determines the dust temperature distribution and, consequently, it influences the shape of far-IR (FIR) continuum dust emission. We sample the compactness using a lognormal Gaussian distribution with a mean of and a standard deviation of 0.4 for each particle (Kapoor et al., 2021; Trčka et al., 2022). The ISM pressure only affects the strength of several emission lines but does not alter the shape of the SED (Groves et al., 2008). Therefore, we treat it as a constant value of . The final one, covering factor of the photodissociation region (PDR) describes the extent that HII regions are coated (Kaufman et al., 1999; Pound & Wolfire, 2022). Since the coating will eventually clear in PDR, this parameter is set as , where represents the age of the stellar particle, and is the molecular cloud clearing or dissipation time scale. means that this star-forming region is thoroughly transparent, while indicates that it is completely covered. We adopt the value of 3 Myr, as recommended in Trčka et al. (2022).
@@
The next step involves modeling the distribution of dust elements within each galaxy. The density of dust at any given volume is determined by the assumption that a predetermined percentage of the metals present in the ISM gas are encapsulated within dust grains, as follows:
@@
where and represents the metallicity and density of the gas, respectively. Also, denotes the dust-to-metal ratio. This ratio is correlated with stellar mass, according to work by Lu et al. (2023), however, we assume it to be a constant of 0.3 for simplicity. Furthermore, we distinguish between the hot circumgalactic medium (CGM) and the cooler ISM gas using two distinct dust recipes:
@@
(i) Only gas cells with non-zero SFR or with a temperature, , below are considered as ISM (Camps et al., 2016):
@@
where is calculated from internal energy, , and electron abundance, , as recommended in TNG website 111https://www.tng-project.org/data/docs/faq/#gen6 as follows:
@@
where , , , and are the adiabatic index, the Bolzmann constant, the hydrogen mass fraction, and the mass of proton respectively. The condition captures gas particles that are eligible for star formation, but were not actually converted into a star-forming region. Meanwhile, the temperature cutoff value of for non-star-forming gas particles ensures that only gas capable of forming dust is considered, as dust cannot form or is rapidly destroyed in hot gas ((Camps et al., 2016; Guhathakurta & Draine, 1989)). The cutoff temperature is arbitrary and empirical. However, this condition does not cause a substantial difference in the dust geometry (see Section 4.1), since the diffuse dust content of each galaxy is essentially determined by the star-forming gas alone (Rodriguez-Gomez et al., 2019; Kapoor et al., 2021).
@@
(ii) The rotationally supported interstellar gas, settled in the disc, are distinguished from the hot CGM gas following Torrey et al. (2012):
@@
where is temperature calculated using Equation 3 and is gas density directly obtained from snapshot data. This condition has been shown to effectively remove cells in the hot halo (Torrey et al., 2012; McKinnon et al., 2016). To enhance the adaptability of GalaxyGenius, we implement both dust recipes, allowing users to freely choose between them. Furthermore, in TNG, hydrodynamics for the gas component is solved using a moving Voronoi mesh via the AREPO code (Weinberger et al., 2020), eliminating the need for smoothing length. The necessary properties for stars, star-forming regions, and dust elements along with their origins are summarized in Table 1.
@@
We employed SKIRT to perform the radiative transfer procedures. SKIRT is a versatile and widely used Monte Carlo (MC) radiative transfer code 222https://skirt.ugent.be/ designed to model the interaction of light with matter in astrophysical systems (Baes et al., 2011; Camps & Baes, 2015). Developed to study processes such as dust absorption, scattering, and thermal re-emission, SKIRT is capable of generating realistic synthetic observations of galaxies, star-forming regions, and other celestial objects. A key feature of SKIRT is its ability to incorporate complex geometries, spatial distributions of stars and dust, and wavelength-dependent properties of materials. Through MC techniques, SKIRT simulates the paths of photon packets as they propagate through a medium, accounting for scattering, absorption, and re-emission processes. SKIRT is particularly well-suited for creating mock observations from hydrodynamical simulations, such as IllustrisTNG and EAGLE (Nelson et al., 2018; Rodriguez-Gomez et al., 2019; Shen et al., 2020; Camps et al., 2016; Guzmán-Ortega et al., 2023; de los Rios et al., 2023; Bottrell et al., 2024; Baes et al., 2024). In such scenario, SKIRT transforms theoretical outputs, such as stellar population distributions and gas content, into observable quantities, including images and SEDs across various wavelength coverage, from UV to FIR. The execution file for SKIRT, with a extension, can be generated through a user-friendly command-line Q&A session.